# Instance-Optimal Imprecise Convex Hull

### v3

@@ Appendix A Further related work
While the strategy is instance-optimal, no explicit retrieval program or running time analysis is provided. Constructing such a program would be a substantial contribution in its own right. Nevertheless, we make two observations that any such retrieval program would need to address. First, the algorithm requires maintaining, at all times, whether any partly or dependent region exists. A natural approach is to dynamically maintain two sets: $A\subset F$, the set of partly regions, and $B\subset F$, the set of dependent regions. However, these sets can exhibit $\Theta(n)$ recourse: after retrieving a single region, up to $\Theta(n)$ other regions may appear on the outer convex hull and become partly. It remains unclear whether these sets can be maintained in amortised constant time. Second, the program must compute, for each selected partly or dependent region $R$, a corresponding witness set $W$. The definition of a witness can be found in [3]. We note that the witness $W$ of a region $R$ cannot be precomputed: the set of regions that may be a witness for $R$ can change dramatically after a retrieval.
@@
The algorithm by Löffler and Raichel [18].
Löffler and Raichel [18] consider the original input set $F$ and partition it into three categories: always regions, never regions, and a third group consisting of partly and dependent regions. Their retrieval strategy is simple: they retrieve all regions in this third group. Let $W(F)=\max_{P\sim F}r(F,P)$ denote the worst-case number of retrievals over all realisations $P\sim F$. Their algorithm retrieves $W(F)$ regions. As such, their result is not instance-optimal but worst-case optimal given $F$ – a strictly weaker guarantee. This makes it easier to achieve, since the set of $W(F)$ regions can be identified during preprocessing.

### v4

@@ Appendix A Related work
While the strategy is instance-optimal, no explicit retrieval program or running time analysis is provided. Constructing such a program would be a substantial contribution in its own right. Nevertheless, we make two observations that any such retrieval program would need to address. First, the algorithm requires maintaining, at all times, whether any partly or dependent region exists. A natural approach is to dynamically maintain two sets: $A\subset F$, the set of partly regions, and $B\subset F$, the set of dependent regions. However, these sets can exhibit $\Theta(n)$ recourse: after retrieving a single region, up to $\Theta(n)$ other regions may appear on the outer convex hull and become partly. It remains unclear whether these sets can be maintained in amortised constant time. Second, the program must compute, for each selected partly or dependent region $R$, a corresponding witness set $W$. The definition of a witness can be found in [bruce2005efficient]. We note that the witness $W$ of a region $R$ cannot be precomputed: the set of regions that may be a witness for $R$ can change dramatically after a retrieval.
@@
The algorithm by Löffler and Raichel [loffler2025preprocessing].
Löffler and Raichel [loffler2025preprocessing] consider the original input set $F$ and partition it into three categories: always regions, never regions, and a third group consisting of partly and dependent regions. Their retrieval strategy is simple: they retrieve all regions in this third group. Let $W(F)=\max_{P\sim F}r(F,P)$ denote the worst-case number of retrievals over all realisations $P\sim F$. Their algorithm retrieves $W(F)$ regions. As such, their result is not instance-optimal but worst-case optimal given $F$ – a strictly weaker guarantee. This makes it easier to achieve, since the set of $W(F)$ regions can be identified during preprocessing.